A 19-inch rope is to be cut into two pieces so that one piece is 3 inches longer than twice the other. How long is each piece, if a = 2, b = 3, and l = 19?
Answer by checkley77(12844)

19=A+B
19=A+(2A+3)
19=3A+3
19-3=3A
3A=16
A=16/3
A=5 1/3
B=2(5 1/3)+3=10 2/3+3=13 2/3
PROOF:
19=(5 1/3)+(13 2/3)=19
19=19